### v1

# An Initial Exploration of Default Images in Text-to-Image Generation

Abstract.
In the creative practice of text-to-image generation (TTI), images are generated from text prompts. However, TTI models are trained to always yield an output, even if the prompt contains unknown terms. In this case, the model may generate what we call default images: images that closely resemble each other across many unrelated prompts. We argue studying default images is valuable for designing better solutions for TTI and prompt engineering. In this paper, we provide the first investigation into default images on Midjourney, a popular image generator. We describe our systematic approach to create input prompts triggering default images, and present the results of our initial experiments and several small-scale ablation studies. We also report on a survey study investigating how default images affect user satisfaction. Our work lays the foundation for understanding default images in TTI and highlights challenges and future research directions.
text-to-image generation, Midjourney, AI art, prompt engineering
copyright: rightsretainedjournalyear: 2025
Figure 1. Examples of default images created with Midjourney, a popular image generator.
[…]
The figure shows different default images, resulting from unrelated input prompts. It is clearly visible that these images have a common theme, motive (i.e., subjects), and style.
1. Introduction
Generative artificial intelligence (AI) has swept into mainstream. One popular type of generative AI is text-to-image generation (TTI), where images are generated by prompting the generative model with text (Oppenlaender et al., 2024; Oppenlaender, 2024; Feng et al., 2023). Most often, the prompt consists of a textual description of the envisioned image, including specific keywords that are known to trigger specific subjects and styles (Oppenlaender, 2023; Placido, 2025). The practice of ‘prompt engineering’ is key to text-to-image generation (Oppenlaender, 2024), where practitioners craft a prompt, observe the outcome, and adjust the prompt until a satisfactory outcome is achieved (Oppenlaender et al., 2024).
While there is a growing number of studies investigating TTI and the practice of prompt engineering, little attention has been paid to what happens when the model does not (in anthropomorphic terms) understand its inputs. TTI models are trained to always yield an output, even if the input prompt contains terms and concepts not known to the model. Therefore, when such a prompt is given to the TTI model, it may generate what we call default images – images that are seemingly unrelated to the input prompt and often do not match the user’s expectations. One key characteristic of default images is that the same (or very similar) default image may appear when prompted with different, unrelated prompts (see Figure 1 and Figure 3).
We argue that studying default images is important for the creative practice of prompt engineering for TTI. When a generative model responds with default images, it is a clear indication that the prompt given to the model is not understood by the model. At the same time, default images reveal gaps in the TTI model’s vocabulary and limits in its capabilities. Consider, for instance, a creative practitioner prompting a generative model with the Tagalog word “pagpag” (referring to a dish made from leftover food scavenged from garbage dumps) and receiving a default image back (e.g., a painting of a woman with a cloud above her face). This is a clear indication that term was not understood and the practitioner should revise the wording of the prompt to improve the generated images. Default images can put an early end to a potentially unfruitful pathway in the practitioner’s prompting of the model. However, default images could also derail the creative process and the quest to generate the desired image. Therefore, default images and their recognition play a crucial role in the creative practice of prompt engineering and for discovering the limits of what a TTI model is capable of. To date, default images have not been investigated in prior research.
In this paper, we present the first preliminary exploration into default images in Midjourney, a popular image generation system. We describe our approach to systematically create a set of input prompts likely triggering default images, and describe our experiments in generating images from these prompts. We then apply affinity diagramming to identify similarities between the generated images, resulting in a set of ten default images. We describe these images and make nine postulates about their characteristics. We further conducted a user study investigating whether default images affect user satisfaction, and how they align with human perception and expectations. Note that due to the vastness of Midjourney’s latent space, this initial investigation is not exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and the importance of studying default images.
The paper makes the following contributions:
  • We define the concept of default images and provide examples of this phenomenon.
  • A description of our systematic approach for generating default images. Due to the vastness of Midjourney’s latent space, this investigation can never be exhaustive. We highlight the challenges of identifying unique but realistic input prompts that reliably trigger default images in Midjourney, and present the results of several ablation studies.
  • A crowdsourced user study investigating how default images affect user satisfaction.
  • We make nine postulates, based on the results of our experiments and study, detailing the characteristics of default images in generative models, and providing a basis for future testing and theoretical development.
2. Related Work
Our review of related work covers three intertwined topics: 1) how text-to-image models are built and prompted, 2) known failure modes and “degenerate” outputs in generative systems, and 3) user-facing evaluations of generation quality and satisfaction.
2.1. Text-to-Image Generation and Prompt Engineering
Text-to-image generation has emerged as a powerful subfield of generative AI, in which deep neural networks (e.g., transformer-conditioned diffusion models) learn to translate free-form textual prompts into high-fidelity images. Prominent systems such as Midjourney, OpenAI’s DALL·E, and the Stable Diffusion family demonstrate the ability to produce photorealistic or artistically stylized images across a wide variety of domains. Generative TTI models are pre-trained on extensive datasets and use probabilistic sampling from their learned distribution to produce images corresponding to the prompt. The knowledge space encoded in the generative model is called latent space – a high-dimensional space learned by the generative model during pre-training.
The creative practice of prompt engineering has emerged as a way to craft and optimize prompts for generative AI systems (Oppenlaender, 2022, 2024). However, Gafni et al. point out that while the generative system attempts to match human expectations, the prompting language is loose and it is near impossible to describe every detail of an envisioned output in just a single textual prompt and with a limited number of characters (Gafni et al., 2022). Further, the model may lack relevant prior knowledge and context. While using special keywords in the prompt can address some of these shortcoming, such as improving the resulting image quality and imbueing it with certain styles (Oppenlaender, 2023), TTI models have specific failure modes that we review in the following section.
2.2. Failure Modes in Generative Models
Generative TTI systems frequently diverge from user intent due to limitations of prompting as an interface to generative AI (Morris, 2024). One reason for this divergence is semantic ambiguity. Polysemous terms – such as “bank”, “bat”, or “ball” – can yield images that reflect unintended senses of a word, creating a mismatch between the user’s mental model and the generative model’s output (White and Cotterell, 2022; Rapp et al., 2025). Users may intentionally use polysemous and ambigous terms in so-called “magic terms” – lexical additives intentionally designed to induce randomness or probe latent failure boundaries – which often result in highly unpredictable or surreal imagery (Oppenlaender, 2023; Maldaner et al., 2025; Quaye et al., 2024). Such phenomena are relevant in interfaces for reviewing and auditing TTI models (Maldaner et al., 2025). Quaye et al., for instance, presented Adversarial Nibbler, a system for crowdsourcing the collection of adversarial examples (Quaye et al., 2024).
Beyond adversarial prompts, training artifacts can manifest as bizarre text-image associations. As text-image pairs have been drawn for the internet, there may be incorrect assignments of text-image pairs (Feng et al., 2023). For instance, rare or out-of-context tokens may surface anomalous content, as evidenced in the analysis by Daras and Dimakis who found “strange” associations between special terms in prompts and concept in generated images (Daras and Dimakis, 2022). They speculated that these associations in TTI models constitute a “secret language”. Williams et al. investigated the effects of non-mainstream dialects in TTI (Williams et al., 2024), finding that minimal changes in prompts (such as dialect) can manifest biases in TTI models. Community reports of special tokens in language models (and by extension, in TTI models), such as “SolidGoldMagikarp”, highlight spontaneous ‘glitch’ phenomena that may arise without adversarial intent (LessWrong, nd; Rumbelow and mwatkins, 2023).
In addition to semantic and training-related failures, TTI models exhibit high stochasticity, introducing a form of variability that itself constitutes a failure mode. Identical prompts can produce substantially different images across runs which complicates both user expectation and scientific reproducibility. Strategies such as sampling multiple random seeds and aggregating outputs have been proposed to mitigate this variability (Liu and Chilton, 2022), and studies show that seed choice alone can influence whether rare or novel concepts emerge (Samuel et al., 2023). Perhaps most critically for creative end-users, failures of prompt recognition – where the model fails to parse novel or unsupported inputs and defaults to placeholder imagery – remain underexamined. Our work targets this gap by presenting a systematic approach for eliciting, identifying, and classifying default-image failures in Midjourney.
2.3. Experiments and User Studies on Generated Image Quality and Satisfaction
Evaluations of text-to-image models often rely on automatic metrics and curated benchmarks, yet these measures are known to misalign with human perception (Kynkäänniemi et al., 2023; Otani et al., 2023). The Fréchet Inception Distance (FID), for instance, can yield scores that disagree with user judgments of image quality and prompt alignment, limiting its utility as a sole evaluation criterion (Kynkäänniemi et al., 2023; Hartwig et al., 2025; Otani et al., 2023). Complementary metrics – such as CLIPScore and VQA-based approaches – have been proposed to capture semantic fidelity, but they too exhibit gaps when compared to subjective assessments (Lin et al., 2024). Recognizing these shortcomings, Russo suggested criteria for a creative TTI benchmark that incorporates insights from linguistics, cognitive psychology, and art theory to better capture human notions of creativity and relevance (Russo, 2022). More recently, the Holistic Evaluation of Text-to-Image Models (HEIM) benchmark systematically assessed 26 state-of-the-art TTI generators across twelve dimensions – ranging from aesthetics and originality to bias and robustness – using a combination of automated measures and human ratings to provide a richer, multidimensional view of model performance (Lee et al., 2023).
Alongside these benchmarks, human‐centered studies bridge the gap between automatic scores and actual user experience. For instance, the Pick-a-Pic dataset gathered over half a million human preference annotations, enabling the training of PickScore, a model that predicts human rankings of generated images with superhuman accuracy, and suggesting a more reliable protocol for evaluating TTI outputs (Kirstain et al., 2023). Building on this, ImageReward constructed a preference‐based reward model from 137,000 expert comparisons and demonstrates that fine-tuning diffusion models against this reward function produces images that users consistently rate as higher quality (Xu et al., 2023). Otani et al. highlight the pervasive lack of verifiable, reproducible human evaluation in TTI research and propose a standardized crowdsourced protocol (Otani et al., 2023). Their results reveal that automatic metrics frequently diverge from human judgments, underscoring the need for rigorous user studies. In interactive settings, tools such as PromptCharm (Wang et al., 2024) have been shown to help participants craft prompts that yield images better aligned with their intent, with users reporting increased satisfaction and perceived control over output quality, while multi-turn guidance systems, such as DialPrompt (Liu et al., 2024), achieve significant gains in both image quality ratings and user‐centricity scores compared to baseline prompt‐engineering interfaces. Finally, exploratory work in art and design teacher education demonstrates how practitioners reflect on prompt strategies and generator behaviors, revealing both learning affordances and cognitive challenges when integrating TTI tools into creative pedagogy (Ringvold et al., 2023).
Together, these experiments and user studies illustrate the complex interplay between objective metrics, human preference, and interaction design in assessing and improving text-to-image generation systems.
3. Default Images in Text-to-Image Generation
3.1. What are Default Images?
Some inputs to text-to-image generation systems may seemingly fail to generate the intended results from the creative practitioner’s perspective. One particular failure mode is what we call default images. The term ‘default’ here is used to denote that TTI models resort to generate these images when they do not recognize their input. This could be, for instance, because the input terms were not part of the generative model’s training dataset or because the terms in the prompt do not invoke strong imagery.
The defining characteristic of default images is that they are encountered in response to multiple different and—crucially—unrelated prompts. The examples given in Figure 1 and Figure 3 demonstrate how different unrelated prompts trigger very similar default images in Midjourney – images with a specific motive and style that originate from a region in the TTI model’s latent space that is not strongly associated with known concepts. Another defining characteristic of default images is that, while there are slight variations, the default images are remarkably stable in terms of their motive (i.e., subject) and style (cf. Figure 1). Their recurring motive and style makes default images distinctly recognizable.
Default images arise primarily due to the way text-to-image models structure their internal representations of visual concepts. When presented with ambiguous or unknown inputs, these models rely on their learned latent space, a multidimensional embedding that encodes visual relationships. Certain regions of this latent space correspond to generic or frequently occurring visual patterns, which the model defaults to in the absence of specific semantic guidance. Additionally, stochastic sampling in image generation contributes to the variability of these default visuals. While overtraining or memorization of frequent training examples can reinforce such biases, the core mechanism behind default images is the inherent structure and learned biases within the model’s latent representations.
A metaphor to think about default images is as follows. We can imagine the latent space as a vast landscape. Images are generated by sampling from regions in this landscape. Known concepts are easy to spot in this landscape (we could think of them as mountains and hills that attract attention). But there are also planes in this landscape that connect vast regions in latent space. These planes do not correspond strongly to specific learned concepts. We like to think that default images result when the TTI models samples from these planes because of the absence of strong ‘attractors’ (known concepts) in the sampled region of latent space.
Due to different training datasets used in pre-training, each TTI model has its own characteristic set of default images. Further, there are differences between default images in different model versions. Due to differences in training materials, it is, for instance, entirely possible that default images in FLUX, an image generation model, may have a more photographic appearance compared Midjourney’s default images, which often tend to be intricate computer graphics and paintings.
3.2. Importance of Studying Default Images in Text-to-Image Generation
Many users on Midjourney type descriptive prompts containing specific keywords, representing known concepts and style modifiers (Oppenlaender, 2023). Further, known concepts in a prompt are like ‘magnets’ that ‘overshadow’ any potentially unknown concepts, shifting the output away from default images to images related to the known concept. This means the majority of Midjourney users will likely not encounter default images in their creative practice of prompt engineering. So why study default images?
Default images provide valuable insights into the limitations and robustness of TTI models. When models generate similar outputs from diverse, unrelated prompts, this indicates semantic gaps and challenges with linguistic constructs, niche topics, or abstract concepts, as well as a lack of generalization. Such patterns suggest that the model may rely on memorized outputs rather than generating novel concepts, revealing limits in the in the model’s computational creativity. Identifying these issues aids in benchmarking model capabilities, benchmarking TTI model performance, and detecting vulnerabilities related to adversarial prompting.
Understanding default images also informs data and model quality improvements. When models default to similar outputs, it may indicate that the training data lacks diversity or contains biased patterns. For instance, it may highlight insufficient coverage of low-resource languages. Addressing these marginal cases can enhance dataset curation and fine-tuning, promoting more nuanced and creative responses.
For creative practitioners and prompt engineers, analyzing default images is valuable when exploring computational creativity and conducting comparative benchmarking of different models. The occurrence of default images during creative prompting reveals the model’s limitations in interpreting abstract or surreal concepts. This understanding helps refine prompt engineering techniques to circumvent “dead zones” where the model consistently produces generic or repetitive outputs. Additionally, identifying when default images appear can inform strategies for recovering from failures, helping artists and designers modify their prompts to yield more meaningful and unique visual outputs.
From a user experience perspective, understanding default images could inform more resilient and user-friendly interface design. When users encounter similar outputs from diverse prompts, it can be frustrating, especially when exploring creative possibilities. Recognizing patterns in default image generation can support the design of better feedback mechanisms, suggesting alternative prompt formulations or guiding users through effective prompt engineering practices. Incorporating failure recovery strategies into TTI systems enhances the user experience by helping users navigate the model’s limitations more effectively.
Default images could also potentially have security implications. An attacker could infer from the appearance of default images the model’s limitations or gaps in training data, which could be exploited to reverse-engineer aspects of the model’s dataset or architecture. Default images can act as a signature of a specific model version or training dataset, which would allow attackers to craft prompts that effectively identify the TTI model or version being used. This information could be valuable for launching more targeted attacks. Knowing that a model produces a predictable default image for certain prompts could allow attackers to inject misleading or confusing input in systems that rely on generative models. As pointed out by Daras and Dimakis (Daras and Dimakis, 2022), TTI systems filter prompts that violate the policy rules, and default image prompts could, potentially, be used to bypass these filters. For instance, in a pipeline that uses TTI models for automated content generation, default images may inadvertently be mistaken for valid outputs.
4. Method
We present the first investigation into default images, acknowledging that in the context of high complexity in TTI models, we can only ever get a tiny glimpse into the latent space and its default images. Our approach consists of systematic prompt creation (Section 4.1), image generation experiments and ablation studies to probe the latent space (Section 4.2), and identification of default images using affinity diagramming (Section 4.3), followed by a survey study (Section 4.4).
4.1. Input Prompts Likely Triggering Default Images
In our initial exploration of the phenomenon of default images, we explored three categories of prompts that could result in a conceptual mismatch between the resulting image and user expectations. The first category includes prompts that are likely to confuse the Midjourney model. This category includes tokens that were likely not part of the model’s training data, such as email addresses, polysemous words, ambiguous topics, and censored words (e.g., political topics). While these prompts are valuable for testing the limits of Midjourney’s models, the probability of default images appearing is too low in our experience. This causes the prompts of this first category to be unsuitable for our systematic approach for generating default images.
The second category involves prompts that are not recognized by the model and lack practical purpose. These prompts include, but are not limited to: contextually misleading phrases (e.g., Einstein’s relativity theology), random character strings and special characters (e.g. aölkjnewjbtn and %), misheard lyrics, and abstract and metaphorical concepts that may not invoke strong imagery (such as absolute ownership and extreme aptitude). While prompts in this category may lead to default images, we excluded them from our analysis due to the limited relevance in real-world applications. These prompts do not reflect how users would typically interact with the TTI model, although they are relevant to for future investigations.
The third category is the most promising to our endeavor, combining both terms not recognized by the model and practical use cases. These prompts are likely used in the real world, but still challenge the model. This includes:
  • A1:
    Rare or uncommon names, such as fantasy names, e.g. Kirk Kirkson
  • A2:
    Corrupted words or altered words, e.g. tractoajrte
  • A3:
    Web addresses, e.g. www.facebook.com and www.quora.com
  • A4:
    Words from low-resource languages, such as A4.1 Finnish, e.g., hevonen (horse), and A4.2 Tagalog (an Austronesian language spoken mainly in the Philippines), e.g., gubat (forest)
  • A5:
    Glitch tokens (Rumbelow and mwatkins, 2023), e.g. rawdownload
  • A6:
    Syllables or abbreviations, e.g. BCD and YMCA
The six categories of prompts emphasize real-world usefulness and they are our focus area and inspiration for creating default images. The low-resource languages were chosen for their language-native words that are not borrowed from other languages and, thus, do not cognate with widely known foreign languages. We focus on Finnish in this study (see Appendix A4.1), leaving Tagalog (Appendix A4.2) to future work. Using this approach, we created a set of prompts, listed in Appendix A. Each of the prompts was checked on haveibeentrained.com, a website that allows to text against the training data used in many TTI models (including LAION (LAION.ai, nd; Schuhmann et al., 2022)). We then generated images with these prompts on Midjourney, as described in the following section.
4.2. Image Generation Experiments and Ablation Studies
Our experiments were executed on Midjourney’s ’Create’ page using fixed parameters listed in Table 1. For each of the six different categories of inputs A1–A6, we ran 20 trials (with exception of category glitch tokens, for which we ran 10 trials due to limited availability of glitch tokens), introducing slight variations between each run by corrupting or altering words (see the prompts in Appendix A). This resulted in 440 images. The images were then visually analyzed, as described in the following section.
Table 1. Image generation parameters used in our experiements
Parameter Value Description
model 6.1 setting the model version
aspect ratio 1 (square) standardizing the format of all generated images
stylize 0 controlling the artistic flair
chaos 0 controlling having surprising random elements
weird 0 controlling having quirky and unconventional elements
seed value 123456 ensuring consistency and reproducibility
As a sensitivity analysis to test the robustness of default images, we conducted three small-scale ablation studies. The first ablation study investigated the effect of different seed values on the default image (see Section 5.2.1). The second study examined default images in different models (see Section 5.2.2), and the last study investigated the default image behavior within larger prompts including known terms (see Section 5.2.3).
4.3. Identification of Default Images
For identifying default images in the set of generated images, we used affinity diagramming. Affinity diagramming (or KJ Method) (Scupin, 1997) is a qualitative method used to organize data into groups based on perceived similarity. The method is commonly employed to categorize complex data. The process, in our case, involved the first two authors collaboratively grouping images on a digital canvas, based on their intuitive understanding of similarity. During the session, the two authors frequently discussed their rationale, iteratively refining the groups. Simultaneously, the two authors developed descriptive labels for the images (e.g., “Lady-Birdhead”), enabling them to refer to these images in their conversation. This collaborative process was continued until all images were either identified as default image (meaning, strong similarities were identified between multiple images created with different prompts) or not. Note that identifying similarities between the images was often straightforward, as depicted in Figure 3.
4.4. Survey Study
In addition to our visual analysis with affinity diagramming, we utilized crowdsourcing on Prolific111www.prolific.com, an online human subject pool, to inquire about people’s perception of default images. In particular, default images may negatively affect user satisfaction when prompting TTI models. To investigate this, we designed an online study with a questionnaire in which four images were presented, including the prompt that created the image. Unknown to the participant, two of the images (Q1 and Q2) had one term in the prompt replaced with a term that caused default images in our experiments. For instance, in Q1, we replaced the term ‘raven’ in the prompt “photoshoot of a woman wearing a white off-shoulder dress with a black raven on top of her head” with the word ’greagoft’. The participant, however, was shown the original prompt and the image resulting from the modified prompt. This setup simulates the surprising mismatch in expectations that an encounter with a default image may produce in practice. For the third question (Q3), we used an image that matches exactly the descriptive prompt, without replacements. In the fourth question (Q4), we used a prompt which is unknown to the model and, thus, the accompanying image is a default image. The four prompts and images are depicted in Figure 6. Participants were asked to rate their satisfaction with each image, given the prompt, on a scale from 1 (Strongly Dissatisfied) to 7 (Strongly Satisfied). The reward to participants was £6.29 per hour, and the median time to complete the survey was 3 min and 49 seconds.
5. Results
5.1. Default Images
Figure 2 showcases the ten default images identified in our study, based on the set of input prompts developed in Section 4.1. Some default images occurred more frequently, as depicted in Figure 4. Figure 3 showcases how different prompts result in the same default image (in this case, ‘Lady-Birdhead’), with only minor variations.
Figure 2. Default images identified in the study. A label for each image is provided to enable future research.
[…]
Figure 3. An example of how different prompts (listed below each image) generate the same default image in Midjourney, with slight variations.
[…]
Figure 4. Frequency of occurrence of default images in our experiments. A1–A6 denote the different test sets, as defined in Section 4.1 and listed in Appendix A
[…]
The default images identified in our study exhibit recurring motifs and artistic styles, reflecting the model’s aesthetic biases when faced with ambiguous or unrecognized prompts. Common motifs include ethereal and symbolic representation with stylized human figures (such as often blended with animal elements or nature, highlighting a recurring theme of human-animal fusion), nature and surreal landscapes (such as birds, trees, and floral patterns) intertwined with human forms or abstract backgrounds, fantasy and mythological themes (such as fantasy art, depicting mythological or otherworldly beings), and symbolism and metaphor (such as crows, masks, and abstract patterns), suggesting a focus on artistic and conceptual representation rather than literal realism. The artistic style of the default images could be described as surrealism and expressionism, painterly and Fine Art aesthetics, romanticism with a dark twist, and abstract realism. These stylistic and thematic patterns indicate that the model, when uncertain, defaults to imaginative and surreal representations, combining human, animal, and natural motifs within a painterly aesthetic framework.
Note that the default images presented in Figure 1 are from a larger set of prompts created by the last author, who has written over 17,000 prompts on Midjourney, many of which were specifically intended to test the limits of the models’ vocabulary and capabilities. These additional images were added in this paper to illustrate the phenomenon of default images, but were not part of this study.
5.2. Ablation Studies
To evaluate the consistency of the default images generated by Midjourney, we examined how altering the seed value, model version, and prompt affected the default images.
5.2.1. Effects of seed values
We visually analyzed how different seed values affect the default images. Using the different seeds with the same prompt resulted in different default images, partially sharing some of the motives (see left images in Figure 5A). While the premise is still the same, one might not associate these images with each other. Thus, the best way to generate default images is by having a fixed seed value for all image generation runs. The images on the right of Figure 5A show how common elements may occur in default images. The commonalities in the default images included portraits of women and animals that were mostly birds.
Figure 5. Ablation studies: (A) Default images generated with different seed values (three images on the left) and image similarities with variations in both seed values and prompts (on the right). (B) The default images on the left demonstrate how known terms (computer, browsing) overpower the unknown terms (in bold), producing images that are not default images. The two images on the right demonstrate how motifs from default images can be created in specific contexts. (C) Default images generated using the prompt “hevonen” and seed 123456 with Midjourney model version 6.1, 6.0, 5.2, and 5.1. Similarities in default images point to similarities in model training data. The four different images generated in response to one prompt (on the left) are the result of Midjourney’s latent space sampling and internal image diversity mechanisms.
[…]
5.2.2. Effects of model versions
We also analyzed how different model versions affect the default images. In our experiments with Midjourney models 6.1 and 6.0, both versions consistently generated the same default images with the same prompt, as depicted in Figure 5C. In contrast, older model versions (e.g. 5.2 and 5.1) generated default images with more surreal aesthetic. These different default images contained the same motifs, such as floating objects, human portraits intertwined with elements of nature, and animals.
5.2.3. Terms triggering default images appearing as part of larger prompts
Last, we experimented with prompts that are known to trigger default images as part of larger prompts to mirror real-world usage, where prompts often contain context or style directions. As shown in Figure 5B, a short prompt with unknown terms may generate default images. However, when those same terms are included in more extensive prompts, the model tends to focus on words that it was able to recognize and generating image from these known words. Using larger prompts resulted in images that no longer shared any characteristics with the default images. The simpler prompt also produces more varied images because it places fewer constraints on the model.
5.3. Default Images and User Satisfaction
5.3.1. Participants
The survey participants ($N=30$) were between 21 and 53 years old (Mean=30.1, SD=6.8) and included 17 men and 13 women. About 53% of the participants had completed a Bachelor degree and 23% a Master’s degree. Participants were familiar with TTI ($Mean=5.8$ and $SD=1.2$ on a scale of 1 (Strongly Disagree) to 7 (Strongly Agree).
5.3.2. User study results
Figure 6 shows that unknown terms potentially triggering default images may negatively affect user satisfaction, but not in all cases. In Q1, the effect of replacing ‘raven’ with ‘greagoft’ resulted in only subtle changes to the resulting image. The image appeared almost as if a raven could be there. The participants’ satisfaction, therefore, was high ($Mean=5,2$, $SD=1,8$). In Q2, where ‘leaves and flowers’ was replaced with ‘greagoft’, the divergence between prompt and resulting image was more noticeable and satisfaction was lower ($Mean=2,7$, $SD=1,9$). A Fisher’s exact test indicates that The way participants perceived the change (subtle vs. noticeable) is significantly related to their satisfaction level ($OR=0.074$, $p<0.05$). The odds ratio of 0.074 indicates that dissatisfaction (1–3 on the Likert scale) is much less likely when the change is subtle compared to noticeable. In other words, participants were significantly more likely to be satisfied (5–7) when the change in the image was subtle (Q1) rather than noticeable (Q2).
For the third question (Q3), we did not switch or omit any words in the prompt, and the image was the exact result of the prompt. Participants were satisfied with the resulting image ($Mean=4,47$, $SD=1,87$). In the fourth question (Q4), we used a prompt known to trigger default images. Participants showed particularly strong dissatisfaction with this image ($Mean=2.6$, $SD=2.0$). Again, there was a significant difference in user satisfaction between Q3 (representing an exact match between prompt and image) and Q4 (respresenting a default image) with $OR=0.21$ and $p<0.05$.
Figure 6. Survey responses, showing how default images can lead to user dissatisfaction.
[…]
6. Discussion
6.1. Default Images in Text-to-image Generation
Default images in TTI generation are an unexplored phenomenon, and we presented the first study on default images. When prompts are vague and feature unknown terms, Midjourney may generate a default image. As prompts become more extensive, with known terms for the model, the model generates images from these known terms and sometimes completely omits unknown terms. When Midjourney does not completely ignore unknown terms, it may generate images sharing visual motifs with motifs in default images. This behavior suggests that the model defaults to familiar visual elements when it encounters words outside its training data. Our findings confirm that these default images are not generated randomly and, on the basis of our experiments, have a narrow visual and thematic range. Although our set included only a tiny sample of default images mainly generated with Midjourney’s version 6.1, we recognized the same motifs across the latest and earlier versions of Midjourney in our experiments. Surrealist elements emerged in common motifs that often depicted human portraits, usually women, and animals, especially birds, along with other abstract botanical or organic forms. Earlier versions of Midjourney have a different set of default images, often including motifs such as balloons, butterflies, floating islands, and intricate paintings (see Figure 1). Default images also have a common style. For instance, in model version 5, the style of some default images could be described as whimsical surrealism with a touch of cartoonish grotesque. Our findings also demonstrate that default images may negatively affect user satisfaction. In the following section, we describe the characteristics of default images.
6.2. Mechanisms and Characteristics of Default Images in Generative Models
Default images reflect inherent biases and patterns encoded in the TTI model. The following nine postulates outline key characteristics of default images, highlighting their occurrence, variability, and evolution as TTI models evolve.
  • P1.
    Emergence of default images due to unknown inputs: TTI models default to certain visual mo (“dtifs and styles (“default images”) when they encounter inputs outside their training data.
  • P2.
    Model-specific default images: Each model version has its own set of default images, based on its training data.
  • P3.
    Paradox of rarity and commonality: Default images are both rare and common. Default images are rare in that they do not occur in response to most prompts written by Midjourney users. However, default images are also common in that they may occur in response to unknown inputs in a prompt. Default images blend infinitely with other known concepts in a prompt.
  • P4.
    Influence of prompt specificity: When users provide vague or ambiguous prompts, default images are more likely to emerge due to the model’s reliance on generic visual patterns.
  • P5.
    Creativity and default image frequency: Default images are more common in ‘creative’ prompts that seek to deviate and innovate.
  • P6.
    Decreasing frequency with model advancement: As the vocabulary of TTI models grows, default images become more rare. Therefore, newer models can be expected to produce fewer default images.
  • P7.
    Modifying default images: It is possible to apply styles to default images.
  • P8.
    Variability of default images per prompt: One prompt can produce several different default images. This is related to how Midjourney leverages stochastic sampling within its generative model to produce diverse visual outputs.
  • P9.
    Evolution of default images with model updates: As Midjourney releases new versions or integrates feedback, the nature and frequency of default images may change, reflecting shifts in the model’s learned visual patterns.
6.3. Explainability and Alignment of Text-to-Image Generation
What does ‘absolute ownership’, ‘nallopand’, or a ‘protruding goofus’ look like? Humans could, after some thinking, come up with some image for these abstract concepts. However, machines struggle to produce a result (cf. Figure 1). This is not surprising, given that these terms were likely not part of the TTI models’ training data. While polysemia and ambiguity will likely continue to be an issue in the alignment of TTI models with human intent (consider, for instance, the prompt “a dog sitting in front of a bank”), default images are a different phenomenon – a type of image that appears in response to prompts that are not understood by the model.
Daras and Dimakis suspected that TTI models have developed a ‘secret language’, and experimented with how gibberish prompts are understood by the model (Daras and Dimakis, 2022). Their hypothesis was that DALL-E 2 had made ‘hidden’ associations between certain terms in text prompts and concepts in images, such as ‘Apoploe vesrreaitais’ meaning ‘bird’. Daras and Dimakis found that these terms often produced consistent results in isolation but also sometimes in combination with other terms in prompts. A similar situation can be found in our work, where some input prompts reliably sample from the same region in latent space. However, unlike in a ‘secret language’, this is not a one-to-one mapping between prompt and resulting image. In our case, several unrelated prompts can produce the same default image, as demonstrated in Figure 3.
Because it is often not clear why the model produced a certain default image, the problem of default images touches on research in the field of interpretability and Explainable AI (XAI). In a sense, the default image – currently – is the mechanism of the TTI system to let the user know that the prompt is not understood. Future work could investigate how a user should be alerted about such prompts, e.g. by designing user interfaces that provide feedback to the user about which parts of the prompt was not understood. Such a solution could help to write better prompts, reducing the appearance of the default images, and thus saving time in the process of prompt engineering. The reduction of both the real and the perceived occurrence of default images would help in producing results that are more satisfying to human perception.
6.4. Limitations and Future Work
We acknowledge limitations in our research. The latent space of TTI models is highly complex, and identifying default images in the context of this complexity is challenging. Midjourney has undergone several version changes since its initial release. Each version of Midjourney produces its own set of default images, and there is a large number of default images to be found in Midjourney’s latent space, with some default images occurring more often than others. Our exploration, therefore, can only give a small glimpse into the vast expanse of Midjourney’s latent space and its default images. On the other hand, given the right inputs, default images are not too difficult to identify, as demonstrated in, for instance, the high frequency of the Lady-Birdhead image (see Figure 4). Future work could investigate default images on a larger scale, for instance, by using technical means – such as feature extraction and clustering, CLIP embedding clustering (Radford et al., 2021), self-supervised learning (e.g., SimCLR (Chen et al., 2020) or BYOL (Grill et al., 2020)), or perceptual and difference hashing – to automatically cluster and identify default images in existing datasets.
We focused on a clearly defined, but limited settings for our experiments. Future work could more broadly study default images with different seed values, parameters, and model versions. In addition, our sensitivity analysis and ablation studies were limited and offer new ground for research. An interesting area for future research that was out of scope in our study is experimentation with other generative models, such as ChatGPT (OpenAI), Stable Diffusion (Stability AI), Imagen (Google), Flux (Black Forest Labs), and Janus (DeepSeek). Another interesting area for future work could be an investigation using Construal Level Theory (Trope and Liberman, 2010), where the more psychologically distant an image is from an individual, the more abstract thoughts it generates. This also relates to research on creativity support in the prompt engineering process, where the effect of default images on the creative process could be investigated.
Having looked at what we call default images, we believe that in the future, this phenomenon will become less frequent due to multiple reasons. First, the available visual material, model vocabulary, and processing power are likely to increase. Second, there will be more efficient models available which will also take into account new factors, such as energy consumption and personalisation. It must be noted, that the phenomenon of default images itself will likely not disappear completely, because we will continue to see situations where a word is not understood by the generative model. A mechanism for handling these situations is needed. Future research could leverage our research to develop tools and interfaces for real-time default‐image detection, and explore techniques to minimize unintended default images.
7. Conclusion
We presented the first systematic investigation on default images in Midjourney. Default images are images with a certain motive and style that occur in response to a variety of different prompts. Through a structured prompt‐generation process and affinity diagramming, we identified ten canonical default images and distilled nine postulates characterizing their properties. Our user study demonstrated that default images may negatively influence user satisfaction. By defining the notion of default images, this work lays the groundwork for future work on the limits of text‐to‐image systems.
Appendix
Appendix A Prompts
A1 – Unique Names: Aranoak Iarnin, Arend Perlaft, Aziel Vartani, Droak Fedril, Fizzah Borandur, Frentl Geddik, Fraoa Seinalur, Jhureth Nyxar, Kaithis Vornex, Meztli Acan, Nuoka Zuberi, Oarn Iznilaker, Papatya Marwood, Pirnak Zerlian, Shelan Creswell, Soraya Lonescu, Torrence Vexley, Tzivani Rhekai, Unnah Nenosim, Uua Lakeiram
A2 – Corrupted Words: Greagoft, Greafirx, Greawarz, Greatird, Greanist, Grealorn, Greamylt, Greavorn, Greaweld, Greatrix, Shuttastfr, Shuttagpdr, Shuttaaips, Shuttagust, Shuttamagu, Shuttagait, Shuttapmt, Shuttagolt, Shuttaiain, Shuttapirn
A3 – URLs: www.amazon.com, www.bing.com, www.chatgpt.com, www.deepseek.com, www.ebay.com, www.facebook.com, www.imdb.com, www.instagram.com, www.linkedin.com, www.nytimes.com, www.paypal.com, www.pinterest.com, www.tiktok.com, www.tumblr.com, www.quora.com, www.reddit.com, www.whatsapp.com, www.wikipedia.org, www.x.com, www.youtube.com
A4.1 – Finnish: mökki, tuoli, pöytä, kasvi, kauppa, maito, kinkku, muovi, mustikka, suola, hevonen, heinasirkka, lapanen, saapas, pelto, lato, käsi, iho, akseli, pultti
A4.2 – Tagalog: apoy aso, bata, daga, dila, galit, gising, gubat, hangin, ilong, iyak, kamay, kulay, lakad, lupa, matanda, ngiti, paa, pagpag, puno, pusa, sapa, takbo, tulog, ulan,
A5 – Glitch Tokens: Rawdownload, PsyNetMessage, UnfocusedRange, Downloadha, RandomRedditorWithNo, InstoreAndOnline, externalActionCode, Embedreportprint, ExternalToEVAOnly, cloneembedreportprint
A6 – Abbreviations: bcd, dwew, fsde, gsod, sfew, gogw, pqwe, fsie, dqaE, gafc, grd, jfdsk, tkrafl, fkeriw, ltr, dfaf, fdfd, rtyu, fgetiu, daqz